For the functions f(x) = 4x^2 - 4
and g(x) = -8x^2 + 4,
find f(g(-3)).

Answer by Nate(3500)

Either:
g(-3) = -8(-3)^2 + 4 = -8(9) + 4 = -72 + 4 = -68
f(-68) = 4(-68)^2 - 4 = 4(4624) - 4 = 18,492
Or:
f(g) = 4(-8x^2 + 4)^2 - 4
f(g) = 4(64x^4 - 64x^2 + 16) - 4
f(g) = 256x^4 - 256x^2 + 64 - 4 = 256x^4 - 256x^2 + 60
f(3) = 256(3)^4 - 256(3)^2 + 60 = 20,736 - 2,304 + 60 = 18,492